OpenStudy (anonymous):

A hydrate of CuSO4 has a mass of 2.82 g before heating. After heating, the mass of the anhydrous compound is found to be 1.80 g. Determine the formula of the hydrate and then write out the name of the hydrate.

OpenStudy (kira_yamato):

Let the formula of the hydrate be CuSO4 . nH2O I'll use the following sets of Ar. Cu = 63.546 S = 32.065 O = 15.9994 H = 1.00794 Mass of anhydrous CuSO4 = 1.80g Thus, no. of moles CuSO4 = 1.80 / (63.546 + 32.065 + 4(15.9994)) = x (Please calculate) No. of moles CuSO4 . nH2O = x Thus, Mr of CuSO4 . nH2O = 2.82 / x 63.546 + 32.065 + 4(15.9994) + n(2(1.00794) + 15.9994) Solve for n.

OpenStudy (kira_yamato):

Values of Ar taken from http://www.ptable.com/